A bullet is fired horizontally with an initial velocity of 865 m/s at a target located 150 m from the rifle.

A) How much time is required for the bullet to reach the target?

B) How far does the bullet fall in this time?
Answer #1

horizontal velocity = V0x = VX = 865 m/sec

horizontal distance = range = x=150 m


a) Range = V0X * t

time = x / vox = 150 / 865 = 0.173 sec


b) distance = y = V0y * t + 1/2 * g * t^2


                    y = 0 + 1/2 * 9.8 * 0.173^2

                        y = 0.29 m